(1) (a) [14.X.15] What is the relationship between the concepts field and force? Wha t are their units? (b) [14X21] An el
Answer #1

(A) force= charge * electric field

F = qE

unit of force is N

unit of electric field is N/C

(b) force = qE

(8*10^-17i - 3.2*10^-16j - 4.8*10^-16k) = -1.6*10^-19*E

E = -500i + 2000j + 3000k

c) force = qE

= 1.6*10^-19 * (2*10^4 i + 2*10^4j)

= 10^-15 *( 3.2i + 3.2j)